OpenStudy (anonymous):

Please help!!!! will fan and give medals!!!!! Solve. x2 – 5x + 6 = 0 A. {3, –2} B. {–2, –3} C. {5, –1} D. {2, 3}

OpenStudy (anonymous):

@acxbox22

OpenStudy (acxbox22):

do you know how to factor?

OpenStudy (paki):

x2 – 5x + 6 = 0 x2-6x+1x+6=0 ..... now make factors and solve...

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

i think it is D.

OpenStudy (paki):

sorry, it will be like this... x2 – 5x + 6 = 0 x2-3x-2x+6=0 x(x-3)-2(x-3)=0 x-3=0 and x-2=0 x=3 and x=2... so you are correct @sammy34daisy